OpenStudy (anonymous):

tan(sin^-1 (5/6) - cos^-1(1/7)) help please.

OpenStudy (freckles):

hey I think I can help you :)

OpenStudy (freckles):

First step: Let a=arcsin(5/6) and let b=arccos(1/7) can you express tan(a-b) so that is in terms of tan(a) and tan(b)

OpenStudy (freckles):

The will be the second step in our process.

OpenStudy (anonymous):

I know the formula is (tan a - tan b)/1- tan a(tan b) and I get ((5/sqrt11) - (4 sqrt3))/1-(5 sqrt11)(4 sqrt3) have no idea ow to simplify.

OpenStudy (freckles):

\[\tan(a-b)=\frac{\frac{\sin(a)}{\cos(a)}-\frac{\sin(b)}{\cos(b)}}{1+\frac{\sin(a)}{\cos(a)} \frac{\sin(b)}{\cos(b)}} \ \cdot \frac{\cos(a)\cos(b)}{\cos(a)\cos(b)} \\ =\frac{\sin(a) \cos(b)-\sin(b)\cos(a)}{\cos(a)\cos(b)+\sin(a)\sin(b)} \]

OpenStudy (freckles):

you are given sin(a)=5/6 and cos(b)=1/7